# Inferring Belief States in Partially-Observable Human-Robot Teams

### v1

@@ VII Conclusion
For decades, the logical predicates model has been a de-facto representation of world states, and has recently been extended to represent belief state prediction for human-robot teaming. The logical structure makes the model compatible with state machines, linear temporal logic planners, and case-based reasoning, which is attractive for real-time systems seeking transparent knowledge representations.
However, the logical predicates model does not consider the probabilistic nature of human cognition – people do not see the world in absolute binary truths, but as probability and managed uncertainty. As such, the logical predicates model is challenged to represent individual human differences in their interpretations of the environment. Moreover, very few works have operationalized the model to human-robot teaming domains, and the literature typically assumes that robot agents have full observability over their environments.
Unable to consider uncertainty, the logical predicates model is limited to low-dimension state spaces and high-certainty predicates. As human-robot teaming domains typically operate in complex environments, misinterpretations can have consequences on downstream planning.
In this work, we explored the strengths and limitations of using a logical predicates belief state model in human-robot teaming. We defined the problem of predicting a co-located human teammate’s belief state, measured by the model’s responses to situation awareness questions. Through a user study we collected a dataset of responses at various points throughout a collaborative cooking game, which we utilized to test the accuracy of a logical predicates model and an LLM at predicting the user’s responses under several degrees of full and partial observability.
@@
    The performance of the logical predicates model only slightly degraded between fully observable and partially observable conditions, supporting the model’s use in more realistic partially observable domains if a careful implementation is feasible.
@@
    The performance of the logical predicates model was matched by a large language model in a zero-shot manner, suggesting that LLMs can be leveraged to help construct logical predicate models, or even supercede them in more complex environments.

### v2

@@ VII Conclusion
For decades, the logical predicates model has been the de-facto representation of world states, and has recently been extended to represent belief state prediction for human-robot teaming. The logical structure makes the model compatible with state machines, linear temporal logic planners, and case-based reasoning, which is attractive for real-time systems seeking transparent knowledge representations.
However, the logical predicates model does not consider the probabilistic and uncertain nature of human cognition. As such, it struggles to represent individual human differences in their interpretations of the environment. Moreover, very few works have operationalized the model to human-robot teaming domains, and the literature typically assumes that robot agents have full observability in their environments. As human-robot teams engage in complex environments, misinterpretation can be consequential to the team’s tasks.
In this work we explored using a logical predicates belief state model in human-robot teaming. We defined the problem of predicting a human teammate’s belief state using the model’s responses to situation awareness questions. Through a user study we collected a dataset of responses at various points throughout a collaborative cooking game, and evaluated a logical predicates model and an LLM at predicting user responses under several degrees of visibility.
@@
    The performance of the logical predicates model did not significantly degrade between fully-observable and partially-observable conditions, supporting the model’s use in more realistic partially-observable domains if a careful implementation is feasible.
@@
    The performance of the logical predicates model was matched by a large language model in a zero-shot manner, suggesting that LLMs can be leveraged to construct or replace logical predicate models.
Figure 7: Starting gameboard of each layout in the user study.